# Interplay of space charge, intrabeam scattering and synchrotron radiation in the Compact Linear Collider damping rings

### v3

@@ IV.1 The resonance
To excite the resonance in this study, a skew sextupolar error is included in the straight section of the lattice. This choice was made in order to avoid any effect from dispersion as present in the arcs, which would create chromatic coupling. Adding only one skew sextupolar error into a non-dispersive region of the lattice will not only excite the resonance but also other (unwanted) skew sextupole resonances. For example, with a single skew sextupole error also the resonance got strongly excited, as revealed by the frequency map analysis for on-momentum particles shown in Fig. 7, where particles are trapped on the resonance showing large diffusion. This study is mainly focused on the resonance and any contribution from other resonances should be avoided.
@@
In order to compensate the strongly excited resonance the Resonance Driving Terms (RDTs) for the and the resonances were computed using PTC from MAD-X [64]. Based on this analysis, a second identical skew sextupole with the same normalized integrated strength was installed towards the end of the same straight section with a phase advance such that the resonance is compensated, while trying to keep the RDT of the resonance the same. The RDT investigation confirms that with the addition of the second skew sextupole, the driving term of the resonance is minimized, while the RDT of the resonance is enhanged. To maintain the same RDT amplitude of the resonance, the skew sextupole strength was reduced by 33%, as shown in Fig. 8.
@@
Following the two skew sextupole configurations, a sensitivity study was performed in order to define a good normalized integrated strength to study the interplay between SC, IBS and SR. The initial beam parameters correspond to the steady state (indicated as SS) and the macro-particle initial conditions are generated to follow a Gaussian distribution in all planes, with 10000 macro-particles that, as shown in Appendix A, is even more than the minimum number needed to have trustworthy results. The working point was set to , which is expected to be significantly affected by the resonances when taking into account the incoherent SC tune shift. The simulation is performed for 14100 turns, which corresponds to the full cycle of the CLIC DRs. Treating the variations of the beam parameters in a self-consistent way can be expensive in terms of computational resources. Thus, in these studies, the frozen SC potential was used. In order to preserve an adiabatic self-consistency of the SC kick during the simulation, the SC potential was recomputed every turns according to the evolution of the transverse beam parameters and the longitudinal line density.
@@
The result of this scan is shown in Fig. 9, where the ratios of the final over the initial values of the horizontal (green) and vertical (red) emittances and intensity (black) are presented as a function of the normalized integrated strength of the skew sextupoles. The horizontal and the vertical emittances are evaluated using the second-order moments of the particle distribution following a similar approach as described in [25]. In particular, a Gaussian fit is applied to the beam profile, and then the weighted RMS emittances are calculated from the actual beam profile up to the extent of the Gaussian fit. For comparison, the vertical emittance evaluated for up to of the distribution is shown in pink. Using the full particle distribution for the computation of the emittance leads to considerable larger emittance values. This is due to the fact that limited number of macro-particles are used in these simulations, outlier particles move to very high amplitudes and in combination with the very large physical aperture of the ring, they have a big impact on the evaluation of the emittance.
@@
Based on this sensitivity study, the integrated strength of the two skew sextupoles was set to in order to have a controllable beam behavior but sufficient blow-up to observe any possible interplay between IBS and SR. For reference, the magnitude of this error corresponds to the case that one of the normal sextupoles of the ring is tilted about the longitudinal axis by about , which is clearly unrealistic. However, in a more realistic scenario, this error could arise from the combination from many of the 400 sextupoles of the CLIC DRs. In this configuration, a vertical emittance blow-up of is observed within a full cycle duration, without any induced particle losses. This very weak response of the beam to this resonance could be related to the fact that the CLIC DRs have a synchrotron motion with a period of about turns, which is quite typical for lepton storage rings, but is fast compared to typical hadron synchrotons. In fact, it was shown that fast tune modulation leads to the excitation of resonance sidebands [65, 66] instead of the usual periodic resonance crossing and the associated particle scattering that is commonly observed in most low-energy hadron rings [67], where the synchrotron motion is multiple times slower.
@@ IV.1.1 Resonance Sidebands
Longitudinal motion modulates the SC induced tune shift that a particle experiences depending on its location in the longitudinal bunch profile. Particles at the tails of the longitudinal bunch profile experience the minimum tune shift while particles located in the core, experience the maximum tune shift. An example of the maximum SC tune shift as a function of the longitudinal position for the vertical tune is shown in Fig. 10. In this example, the tune modulation from chromaticity is not taken into account.
@@
On-momentum particles, located in the core of the longitudinal distribution, do not experience any tune modulation due to the lack of longitudinal motion. In this case, the particles sample the usual betatron resonances of the lattice as shown in Fig. 11b. The initial transverse coordinates of these particles are shown in Fig. 11a. Off-momentum particles, on the other hand, exhibit synchrotron motion that modulates the tune shift they experience due to SC. As particles move closer to the core of the beam, their SC tune shift grows, while as they move to the edges of the longitudinal bunch profile their tune shift is minimized. Each particle passes through the maximum tune shift twice per synchrotron period and thus, the SC tune shift is modulated at even multiples of the synchrotron tune.
@@
The movement of the off-momentum particles in the tune space creates zones of chaotic motion around resonances. The size of the chaotic zones depends on the strength of the resonance and on the frequency of the synchrotron oscillations. Since the SC tune shift is modulated at even multiples of the synchrotron frequency, when synchrotron motion becomes faster, resonance sidebands start to rise. In the case of the resonance the sideband condition will be with an integer number. Figure 12a shows the position of the particles of the initial transverse distribution, where the color-code indicates the tune diffusion over a time span of 5 synchrotron periods, for off-momentum particles at , while Fig. 12b depicts the tune footprint of the same distribution. It becomes clear that the sidebands of the resonance, appearing at even multiples of the synchrotron tune, cause tune diffusion. Another representation is shown in Fig. 13, where the vertical tune as a function of the initial vertical position for particles at is shown and the tune diffusion is again indicated by the color code. Resonance sidebands produce significantly weaker particle diffusion than the usual scattering on the resonance islands, which is observed in most hadron machines where particles have slower synchrotron motion (with synchrotron periods on the order of 1000 turns) [67].
@@
For comparison, a study case was investigated with a different synchrotron period. To change the synchrotron period of the machine, the harmonic number and the RF voltage were modified, keeping the momentum spread constant. Since the properties of the RF bucket are changed, the same momentum spread corresponds to a different bunch length. Since SC strongly depends on the longitudinal line density, to maintain the same SC tune shift as the nominal case, the beam intensity was also modified accordingly. The full list of parameters, for the different synchrotron periods that are used for these studies, are summarized in Table 3.
@@
For this comparison, the synchrotron period was modified to turns, 10 times slower than the nominal case. This case is expected to be in the usual particle trapping regime, with significantly stronger response to resonances. The initial distribution and the tune footprint are shown in Fig. 14a and in Fig. 14b respectively, for off-momentum particles at and with the tune diffusion evaluated over 2 synchrotron periods. It is evident that in this case there are no resonance sidebands excited, but there is a chaotic zone around the resonance due to the synchrotron motion of the particles that leads to variations of their SC tuneshift and thus, to resonance crossing and associated scattering at the resonance islands [67].
IV.1.2 Varying the Synchrotron Period
To investigate the sensitivity to the resonance, a scan on a wide range of synchrotron periods was performed, for the working point , . The synchrotron period was modified as described in the previous paragraph. The initial parameters are, as before, the ones of the steady state. Even though it takes only about 7000 turns to reach the steady state, in this case the duration of the simulations is twice as long, i.e. 14100 turns, to have a clear evolution of the emittances. The skew sextupoles normalized integrated strength is set at . The ratios of the horizontal emittance (green), vertical emittance (red) and intensity (black), for the CLIC DRs operating with different synchrotron periods , are summarized in Fig. 15. For very small synchrotron periods , there is no response of the beam to the resonance. In the range between 1000 turns and 2000 turns the response is maximized while for larger periods the response decreases due to the limited number of resonance crossings, as the synchrotron period becomes comparable to the simulation time of 14100 turns.
A full vertical tunescan around the resonance was performed for four chosen synchrotron periods; , , (nominal) and turns, for a range of vertical tunes and for a constant horizontal tune . All four cases are shown in Figs. 16, respectively. As expected, the case with the slower synchrotron motion shows the largest vertical emittance blow-up of the order of and of particle losses, followed by the nominal case with a blow-up of and of particle losses. In both cases, the losses are reduced when the emittance is blown-up, since the strength of the SC force is also reduced. For the other two faster synchrotron motion cases, the effect of the strongly excited resonance becomes negligible. So clearly, for the nominal synchrotron motion of the CLIC DRs the beam degradation due to SC induced tune modulation close to a resonance is largely suppressed.
@@ V Interplay of SC and IBS
As a first test, simulations were performed without the skew sextupolar error, i.e. the resonance not being excited, to verify that without the effect of the resonance, there is no interplay between the two effects. The working point was set at , , which is of the most affected working points in the previously shown SC studies. Figure 17 shows the vertical emittance evolution when only SC (green), only IBS (blue) and both IBS with SC (red) act on the beam. The error-bars indicate the standard deviation over three different runs, which appear to be negligible in this case. As expected, when only SC is considered in the simulations, the vertical emittance remains unaffected while in the case of the combined simulation with IBS and SC, only the emittance growth as expected from the pure IBS effect is observed.
A full vertical tunescan around the resonance was performed for the nominal synchrotron motion case turns, for a range of vertical tunes and for a constant horizontal tune . This time the skew sextupolar error with a normalized integrated strength , is included, exciting the resonance. Figure 18a shows the relative transverse emittances and intensity when both IBS and SC act on the particle distribution. Results show the sensitivity to the is strongly enhanced when both effects are present, producing a significantly larger vertical emittance blow-up and more losses for the working points affected by the resonance, while IBS is dominant for the working points away from the resonance.
For comparison, the same tunescan was performed for the slow synchrotron period case turns, with the resonance excited. The results are shown in Fig. 18b where a similar behavior as the previous case is observed, with increased vertical emittance blow-up and particle losses.
@@ VI Interplay of SC, IBS and SR
The final step is the addition of the SR effect into the simulations to study the interplay of the three effects in the vicinity of an excited resonance. A tunescan was performed in a range of vertical tunes from 10.32 to 10.38, for the nominal synchrotron motion. Similar as before, the simulations are performed for a beam which is at the equilibrium state. Figure 19 shows the ratio of the final over initial horizontal (green) and vertical (red) emittance and intensity (black). The strong damping from SR strongly mitigates the impact of the other two effects, leaving a maximum residual vertical emittance growth of the order of . Particle losses are significantly less than and are observed only for the working points that were affected the most by the resonance, as shown in Sections IV and V.
@@ VI.1 Full Cycle Simulations
Figure 20 shows the comparison of the tracking simulation including the effects of SC, IBS and SR in blue and the analytical calculations of IBS and SR in red. The equilibrium emittances emerging from SR and QE, as calculated from the radiation integrals, are shown in the dashed green lines. As described earlier, the beam is injected with large transverse emittances. At that point, SC and IBS are expected to be very weak with only radiation damping acting on the beam. Due to the large amplitudes, detuning with amplitude due to the strong sextupoles might also affect the beam, leading to resonance crossing. As the beam becomes smaller, the effects of IBS and SC become stronger leading to larger emittance values than the equilibrium defined by SR and QE, known as steady state of the beam. The tracking simulation with SC, IBS and SR seems to be in excellent agreement with the analytical predictions with only IBS and SR, leading to the conclusion that for the ideal lattice, i.e. without imperfections, SC does not change the final steady state of the operational scenario despite the very small emittances and the large vertical tune spread that SC might induce.
@@
To investigate how the resonance can affect the full cycle of the CLIC DRs, simulations including the skew sextupole errors with normalized integrated strength , were performed. The beam was again initialized to the injected beam parameters and the duration of the simulations is 14100 turns. The horizontal tune is set to the nominal . In particular, to understand the dependence of the interplay effect along the cycle on the working point, the simulations were performed for three different vertical tunes: , and . The results of these simulations are summarized in Figs. 21a and 21b for the horizontal and vertical planes, respectively. In order to highlight the working point dependence, both the horizontal and vertical emittances are normalized to the ones from analytical calculations of IBS and SR, where no errors and no losses are observed.
@@
Comparing the results from the three different working points, it becomes clear that even in the presence of strong SR, the working point choice is very important. For the vertical tune , particles are away from the resonance and no beam degradation is observed. On the other hand, for the other two working points that are closer to the resonance, a strong enhancement both of particle losses and emittance blow-up is observed. In particular, for the , of the particles are lost soon after injection, explaining the fact that the emittance in the simulations is smaller than the one from the analytical calculations during the first part of the cycle (up to around 4000 turns). As the beam is shrinking fast due to the SR effect, the losses are eliminated and the interplay between the IBS and SC effects in the close vicinity of the strongly excited resonance, leads to a final vertical emittance that is higher than the expected one. The horizontal emittance increase observed between 2000 and 4000 turns, following a similar pattern as in the emittance evolution observed in the vertical plane, indicates some interaction of the beam with the coupling resonance, leading to emittance exchange in the two planes. The large fluctuation of the emittance values observed at the beginning of the cycle are associated to the strong particle losses during the first parts of the cycle, where the beam is still large. Fig. 21c shows the particle losses along the cycle for the three working points.
@@
As already discussed earlier, the preservation of the small vertical emittance in damping rings is of great importance. Our studies demonstrated that the working point choice is very important for minimising the impact of the interplay between SC and IBS, even in the presence of strong SR. In a next step, we investigate the sensitivity to the skew sextupolar resonance and define a threshold strength at which the extraction requirements are not met in the case the chosen operational working point is strongly affected by this resonance (i.e. ). To this end, simulations for the full cycle were repeated for different strengths of the skew sextupolar error and the results are summarized in Fig. 22 and Fig. 23, for the vertical emittance evolution and particle losses along the cycle, respectively.
@@
In terms of the vertical emittance, the extraction requirement (red dashed line) is met for all skew sextupole errors shown. For the intensity requirements, there is a degradation even for weak skew sextupoles with a threshold normalized integrated strength between and . However, it should be emphasized that these skew sextupole errors are quite large and far from realistically expected field errors. For comparison, a normal sextupole used for the chromaticity correction in the CLIC DR lattice has a strength of .
@@
In Section IV, it was shown that for slower synchrotron motion, the response to the resonance is enhanced. To investigate how different the performance of the CLIC DRs would be in this case, similar tracking simulations were performed for the working point , for a synchrotron period of turns, and the two weakest errors of the ones presented above were used. Figure 24 shows the comparison of the vertical emittance (a) and intensity (b) ratios, for the nominal synchrotron motion case (full lines) and the slow synchrotron motion case (dashed lines) for normalized integrated strengths of (blue) and (orange). Significantly higher losses are observed in the case of slower synchrotron motion, as already observed in section V. Thus, the synchrotron motion in the CLIC DRs helps mitigating the impact of strong resonances on beam quality, in terms of particle losses and beam emittances.
@@
Another requirement that the downstream machine following the DRs might have is the shape of the beam, as this can have an impact on the luminosity. Excited resonances and IBS can eventually change the shape of the beam profiles by populating the tails more, especially in the longitudinal plane. On the other hand, SR shrinks the profiles back to Gaussian or even narrower [6, 4]. To verify if in this case the profiles remain Gaussian, the initial and the final longitudinal beam profiles are compared with Gaussian profiles for the nominal synchrotron period case and the skew sextupolar error with normalized integrated strength of , which was the worst behaving case. Figure 25 shows a comparison between the initial (red) and final (blue) beam profiles as produced from the simulations (crosses) and are compared with Gaussian fits (solid lines). As indicated, during the full cycle of the CLIC DRs, in the presence of SC, IBS and SR with the resonance excited, the longitudinal beam shape remains Gaussian.

### v4

@@ IV.1 The resonance
To excite the resonance in this study, a skew sextupole error is included in the straight section of the lattice. This choice was made in order to avoid any effect from the dispersion that is present in the arcs, which would create chromatic coupling. Adding only one skew sextupole error in a non-dispersive region of the lattice will not only excite the resonance but also other (unwanted) skew sextupole resonances. For example, with a single skew sextupole error, the resonance gets strongly excited, as revealed by the frequency map analysis for on-momentum particles shown in Fig. 7. Particles get trapped on the resonance as revealed from the large tune diffusion. This study is mainly focused on the resonance and any contribution from other resonances should be avoided.
@@
In order to compensate the strongly excited resonance the Resonance Driving Terms (RDTs) for the and the resonances were computed using PTC in MAD-X [64]. Based on this analysis, a second identical skew sextupole with the same normalized integrated strength was installed towards the end of the same straight section with such a phase advance that the resonance is suppressed. To maintain the same level of excitation in terms of RDT amplitude for the resonance, the strength of the two skew sextupoles had to be reduced by 33% compared to the case of using an individual magnet. The RDT calculations before and after these lattice manipulations are shown in Fig. 8.
@@
With the two skew sextupole configurations, a sensitivity study was performed in order to define a normalized integrated skew sextupole strength that allows studying the interplay between SC, IBS and SR. The initial beam parameters correspond to the steady state (indicated as SS) emittances and the initial particle coordinates are generated to follow a Gaussian distribution in all three planes, with 10000 macro-particles (this choice is based on convergence studies as shown in Appendix A). The working point was set to , which is expected to be significantly affected by the resonances, as shown in Fig. 7. The simulation is performed for 14100 turns, which corresponds to the full cycle of the CLIC DRs. Treating the variations of the beam parameters in a self-consistent way can be expensive in terms of computational resources. Thus, in these studies, the frozen SC potential was used. In order to preserve an adiabatic self-consistency of the SC kick during the simulation, the SC potential was recomputed every turns according to the evolution of the transverse beam parameters and the longitudinal line density of the macro-particle distribution.
@@
The result of a scan of the skew sextupole strength is shown in Fig. 9, where the ratios of the final over the initial values of the horizontal (green) and vertical (red) emittances and intensity (black) are presented as a function of the normalized integrated strength of the skew sextupoles. The horizontal and the vertical emittances are evaluated using the second-order moments of the particle distribution following a similar approach as described in [25]. In particular, a Gaussian fit is initially applied to the beam profile, and the weighted RMS emittances are calculated using the profile up to the extent of the Gaussian fit. For comparison, the vertical emittance evaluated for up to of the Gaussian fitted distribution is shown in pink. Using the full profile for the computation of the emittance leads to considerable larger values. This is due to the fact that a limited number of outlier macro-particles move to very high amplitudes, without getting lost as the ring has a very large physical aperture, and they have a significant impact on the emittance evaluation.
@@
Based on this sensitivity study, the integrated strength of the two skew sextupoles was set to , as it allows for a controllable beam behavior but sufficient blow-up to observe any possible interplay between IBS and SR. For reference, the magnitude of this error corresponds to the case that one of the normal sextupoles of the ring is tilted about the longitudinal axis by about , which is clearly unrealistic. However, in a more realistic scenario, this error could arise from the combination from many of the 400 sextupoles of the CLIC DRs. In this configuration, a vertical emittance blow-up of is observed within a full cycle duration, without any induced particle losses. This very weak response of the beam to this resonance could be related to the fact that the CLIC DRs have a synchrotron motion with a period of about turns, which is quite typical for lepton storage rings, but is fast compared to typical hadron synchrotons. In fact, it was shown that fast tune modulation leads to the excitation of resonance sidebands [65, 66] instead of the usual periodic resonance crossing and the associated particle scattering that is commonly observed in most low-energy hadron rings [67], where the synchrotron motion is multiple times slower.
@@ IV.1.1 Resonance Sidebands
Longitudinal motion modulates the SC induced tune shift that a particle experiences depending on its location in the longitudinal bunch profile. Particles at the tails of the longitudinal bunch profile experience the minimum tune shift while particles located in the core, experience the maximum tune shift. An example of the maximum SC tune shift as a function of the longitudinal position for the vertical tune is shown in Fig. 10. In this example, the tune modulation from chromaticity is not taken into account, and the calculation is done for a particle at the closed orbit.
@@
On-momentum particles, located in the core of the longitudinal distribution, do not experience any tune modulation due to the lack of longitudinal motion. In this case, the particles sample the usual betatron resonances of the lattice as shown in Fig. 11b. The initial transverse coordinates of these particles are shown in Fig. 11a. On the other hand, the SC induced tune shift is modulated for off-momentum particles due to their synchrotron motion. As particles move closer to the core of the beam, their SC tune shift grows, while as they move to the edges of the longitudinal bunch profile their tune shift is minimized. Each particle passes through the maximum tune shift twice per synchrotron period and thus, the SC tune shift is modulated at even multiples of the synchrotron tune.
@@
The movement of the off-momentum particles in the tune space creates chaotic motion around resonances. The size of the chaotic phase space area depends on the strength of the resonance and on the frequency of the synchrotron oscillations. In the case of fast synchrotron motion, resonance sidebands start to arise at even multiples of the synchrotron tune since the SC tune shift is modulated with twice the synchrotron tune. In the case of the resonance the sideband condition will be with an integer number. Figure 12a shows the position of the particles of the initial transverse distribution for off-momentum particles at , while Fig. 12b depicts the tune footprint of the same distribution. It becomes clear that the sidebands of the resonance cause tune diffusion. Another representation of the same tracking study is illustrated in Fig. 13, where the vertical tune as a function of the initial vertical position for particles at is shown. Resonance sidebands produce significantly weaker diffusion than the usual scattering on the resonance islands, as shown in more detail below.
@@
For comparison, a study case was investigated with a different synchrotron period. To change the synchrotron period of the machine, the harmonic number and the RF voltage were modified, keeping the momentum spread constant. Since the properties of the RF bucket are changed, the same momentum spread corresponds to a different bunch length. Since SC strongly depends on the longitudinal line density, the beam intensity was also modified accordingly in order to to maintain the same SC tune shift as the nominal case. The full list of parameters, for the different synchrotron periods that are used for these studies are summarized in Table 3.
@@
For this comparison, the synchrotron period was modified to turns, 10 times slower than the nominal case. This case is expected to be in the usual resonance trapping regime, with significantly stronger beam degradation due to resonances. The initial distribution and the tune footprint are shown in Fig. 14a and in Fig. 14b respectively, for off-momentum particles at and with the tune diffusion evaluated over 2 synchrotron periods. It is evident that in this case there are no resonance sidebands excited, but there is a chaotic area around the resonance due to the synchrotron motion of the particles that leads to variations of their SC tune shift and thus, to resonance crossing and associated scattering at the resonance islands [67].
IV.1.2 Sensitivity to the Synchrotron Period
To investigate the sensitivity to the resonance, a scan on a wide range of synchrotron periods was performed, for the working point , . The synchrotron period was modified as described in the previous paragraph. The initial beam parameters are, as before, the ones of the steady state. Even though it takes only about 7000 turns to reach the steady state, the duration of the simulations is twice as long (i.e. 14100 turns) to observe a clear evolution of the emittances. The skew sextupoles normalized integrated strength is set to . The ratios (final over initial) of the horizontal emittance (green), vertical emittance (red) and intensity (black) for the CLIC DRs operating with different synchrotron periods are summarized in Fig. 15. For very small synchrotron periods (below 100 turns), there is no response of the beam to the resonance. In the range between 1000 turns and 2000 turns the response is maximized while for larger periods the response decreases due to the limited number of resonance crossings, as the synchrotron period becomes comparable to (or in some cases even larger than) the number of turns simulated (i.e. 14100 turns).
A full vertical tunescan around the resonance was performed for four chosen synchrotron periods; , , (nominal) and turns, for a range of vertical tunes , for a constant horizontal tune and for a duration of 14100 turns. All four cases are shown in Fig. 16, respectively. The case with the slower synchrotron motion shows the largest vertical emittance blow-up of the order of and of particle losses, followed by the nominal case with a blow-up of and of particle losses. In both cases, the losses are reduced when the emittance is blown-up, since the strength of the SC force is also reduced. For the two cases with faster synchrotron motion, the effect of the strongly excited resonance becomes negligible. For the nominal synchrotron motion of the CLIC DRs the beam degradation due to SC induced tune modulation close to a resonance is largely suppressed.
@@ V Interplay of SC and IBS
As a first test, simulations were performed without the skew sextupolar error, i.e. the resonance not excited, to identify if there is any interplay between the two effects. The working point was set to , , which is the one with the higher sensitivity to the resonance as shown before. Figure 17 shows the vertical emittance evolution when only SC (green), only IBS (blue) and both IBS with SC (red) act on the beam. The error-bars indicate the standard deviation over three different simulation runs, which appear to be negligible in this case. As expected, when only SC is considered in the simulations, the vertical emittance remains constant while in the case of the combined simulation with IBS and SC, the emittance evolution is dominated by IBS as the growth is the same as the one expected from the pure IBS effect.
A full vertical tunescan around the resonance was performed for the nominal synchrotron motion case turns, for a range of vertical tunes and for a constant horizontal tune . Here the resonance is excited by a skew sextupolar error with a normalized integrated strength . Figure 18a shows the relative transverse emittances and intensity when both IBS and SC act on the particle distribution. The sensitivity to the is strongly enhanced when both effects are present, producing a significantly larger vertical emittance blow-up and more losses for the working points affected by the resonance, while IBS dominates the emittance growth for the working points sufficiently far away from the resonance.
For comparison, the same tunescan was performed for the slow synchrotron period case turns. The results are shown in Fig. 18b where a similar behavior as the previous case is observed, with enhanced vertical emittance growth and particle losses.
@@ VI Interplay of SC, IBS and SR
The final step is including the SR effect into the simulations to study the interplay of the three effects in the vicinity of an excited resonance. A tunescan was performed in a range of vertical tunes from 10.32 to 10.38, for the nominal synchrotron motion. Similar as before, the simulations are performed for a beam at the equilibrium state. Figure 19 shows the ratio of the final over initial horizontal (green) and vertical (red) emittance and intensity (black). The strong damping from SR strongly mitigates the impact of the other two effects, leaving a maximum residual vertical emittance growth of the order of . Particle losses are significantly less than and are observed only for the working points that were affected the most by the resonance, as shown in Sections IV and V.
@@ VI.1 Full Cycle Simulations
Figure 20 shows the comparison of the tracking simulation including the effects of SC, IBS and SR in blue and the analytical calculations of IBS and SR in red. The equilibrium emittances emerging from SR and QE, as calculated from the radiation integrals, are shown in the dashed green lines. As described earlier, the beam is injected with large transverse emittances. At that point, SC and IBS are very weak, and mostly radiation damping is acting on the beam. Due to the large amplitudes, detuning with amplitude due to the strong sextupoles might also affect the beam, leading to resonance crossing. As the beam becomes smaller, the effects of IBS and SC become stronger leading to larger emittance values than the equilibrium defined by SR and QE, known as steady state of the beam. The tracking simulation with SC, IBS and SR are in excellent agreement with the analytical predictions with only IBS and SR, leading to the conclusion that for the ideal lattice, i.e. without imperfections, SC does not change the final steady state of the operational scenario despite the very small emittances and the large vertical tune spread that SC might induce.
@@
To investigate how the resonance can affect the beam during the full cycle of the CLIC DRs, simulations including the skew sextupole errors with normalized integrated strength , were performed. The beam was again initialized to the injected beam parameters and the duration of the simulations is 14100 turns. The horizontal tune is set to the nominal . In particular, to understand the dependence of the interplay effect on the working point along the cycle, the simulations were performed for three different vertical tunes: , and . The results of these simulations are summarized in Figs. 21a and 21b for the horizontal and vertical planes, respectively. In order to highlight the working point dependence, both the horizontal and vertical emittances are normalized to the ones from analytical calculations of IBS and SR, where no errors and no losses are observed.
@@
Comparing the results from the three different working points, it becomes clear that even in the presence of strong SR, the working point choice is very important to avoid beam degradation. For the vertical tune , particles are away from the resonance and no beam degradation is observed. On the other hand, for the other two working points that are closer to the resonance, strong particle losses and significant emittance blow-up are observed. In particular, for , of the particles are lost soon after injection, explaining the fact that the emittance in the simulations is smaller than the one from the analytical calculations during the first part of the cycle (up to around 4000 turns). As the beam is shrinking fast due to the SR effect, the losses are eliminated and the interplay between the IBS and SC effects in the close vicinity of the strongly excited resonance, leads to a final vertical emittance that is higher than the expected one. The horizontal emittance increase observed between 2000 and 4000 turns, following a similar pattern as in the emittance evolution observed in the vertical plane, indicates some interaction of the beam with the coupling resonance, leading to emittance exchange in the two planes. The large fluctuation of the emittance values observed at the beginning of the cycle is associated to the strong particle losses during the first parts of the cycle, where the beam is still large. Figure 21c shows the particle losses along the cycle for the three working points.
@@
Our studies demonstrated that the working point choice is very important for minimising the impact of SC and IBS, even in the presence of strong SR. As a next step, we investigate the sensitivity to the skew sextupolar resonance and define a threshold strength at which the extraction requirements are not met in the case the chosen operational working point is affected by this resonance (i.e. ). To this end, simulations for the full cycle were repeated for different strengths of the skew sextupolar error and the results are summarized in Fig. 22 and Fig. 23, for the vertical emittance evolution and particle losses along the cycle, respectively.
@@
In terms of the vertical emittance, the extraction requirement (red dashed line) is met for all skew sextupole errors examined. On the other hand, losses are observed even for weak skew sextupoles with a threshold normalized integrated strength between and . However, it should be emphasized that these skew sextupole errors are quite large and far from realistically expected field errors. For comparison, a normal sextupole used for the chromaticity correction in the CLIC DR lattice has a strength of .
@@
In Section IV, it was shown that for slower synchrotron motion, the beam response to the resonance is enhanced. To investigate the impact of the synchrotron period on the performance of the CLIC DR, similar tracking simulations were performed for the working point , for a synchrotron period of turns. Figure 24 shows the comparison of the vertical emittance (a) and intensity (b) ratios, for the nominal synchrotron motion case (full lines) and the slow synchrotron motion case (dashed lines) for normalized integrated strengths of (blue) and (orange). Significantly higher losses are observed in the case of slower synchrotron motion, as already observed in section V. Thus, the synchrotron motion in the CLIC DRs helps mitigating the impact of strong resonances on beam quality, in terms of particle losses and beam emittances.
@@
Finally, the profile shape of the beam distribution also need to be investigated, as non-Gaussian profiles can impact the luminosity. Excited resonances and IBS can eventually change the shape of the beam profiles by populating the tails more, especially in the longitudinal plane. On the other hand, SR shrinks the profiles back to Gaussian or in some cases even to profiles with less than Gaussian tails [6, 4]. To verify if in the case of the CLIC DRs the profiles remain Gaussian, the initial and the final longitudinal beam profiles are compared with Gaussian profiles for the nominal synchrotron period case and the skew sextupolar error with normalized integrated strength of , which was the worst behaving case. Figure 25 shows a comparison between the initial (red) and final (blue) beam profiles as produced from the simulations (crosses) and are compared with Gaussian fits (solid lines). As indicated, during the full cycle of the CLIC DRs, in the presence of SC, IBS and SR with the resonance excited, the longitudinal beam shape remains Gaussian.